when solving linear equations with 3 variables, how do you know which one is the easiest to solve for?
Answer by pushpaharan(47)

it's all depends on your judgement.
if one variable in has same co-efficent in different qquations get rid of that first by add/ subtract...
if none.. then look for the variable where co-effients are mutiple of other... sp by mutiplying bt the factor you can make coeffiecnts same
e.g: 4x-3y+2z =10
5x-7y-4z=8
6x+4y-5z= 12
then get rid of z in 1st and 2nd... then 1st and 3rd...